Mary can work word problems twice as fast as Tara. If they work on a mathematics assignment as a team, it takes two hours to finish the assignment. How long would it have taken Mary alone to do the assignment?

Answer by richwmiller(17219)

2/m+2/2m=1
m=3 hours for mary
2m=6
check
2/6+2/3=1
1/3+2/3=1